Please simplify each trigonometric expression:
1. cos 55º cos 25º + sin 55º sin 25º
2. sin 70º cos 10º – cos 70º sin 10º

Answer by lwsshak3(11628)

Please simplify each trigonometric expression:
1. cos 55º cos 25º + sin 55º sin 25º
2. sin 70º cos 10º – cos 70º sin 10º
..
Use following two identities to solve:
cos(s-t)=cos s cos t + sin s sin t
sin(s-t)=sin s cos t - cos s sin t
..
1. cos 55 cos 25 + sin 55 sin 25=cos(55-25)=cos(30)=√3/2
2. sin 70 cos 10 – cos 7 sin 10º=sin(70-10)=sin(60)=√3/2